# A Survey on 3D Gaussian Splatting

### v8

@@ 5.4. Avatar
3D GS has emerged as a powerful tool for human avatar reconstruction, primarily advancing along two directions: full-body modeling and head-centric modeling. For full-body avatars (Jena et al., 2023; Ye et al., 2023; Li et al., 2024k; Lei et al., 2024a; Moreau et al., 2024; Kocabas et al., 2024; Abdal et al., 2024; Zheng et al., 2024b; Hu and Liu, 2024; Jiang et al., 2024a; Hu et al., 2024c; Pang et al., 2024; Qian et al., 2024b; Yuan et al., 2024; Jung et al., 2023; Li et al., 2023b, 2024h), the current methods typically anchor 3D Gaussians in a canonical space and deform them via parametric body models (e.g., SMPL) or cage-based rigging to model dynamic motions. These approaches adopt a hybrid deformation strategy: linear blend skinning handles rigid skeletal transformations such as joint rotations, while pose-conditioned deformation fields account for secondary non-rigid effects like muscle jiggles. For head avatars (Qian et al., 2024a; Dhamo et al., 2023; Xiang et al., 2023; Saito et al., 2024; Chen et al., 2023b; Zhao et al., 2024a; Zhou et al., 2024d; Li et al., 2024j; Rivero et al., 2024), the emphasis shifts to modeling intricate facial expressions, fine-grained geometry (e.g., wrinkles, hair (Luo et al., 2024)), and dynamic speech-driven animations. Techniques mainly combine parametric morphable face models (e.g., FLAME) with deformable 3D Gaussians, employing diffusion strategies and expression-aware deformation fields to disentangle rigid head poses from non-rigid facial movements. Both directions exploit the speed advantage and editability of GS to enable fast training, real-time rendering, and precise control over deformations, while addressing challenges in cross-frame correspondence, topology flexibility, and multi-view consistency.
@@ 5.5. Endoscopic Scene Reconstruction
Endoscopic Scene reconstruction is pivotal in robot-assisted minimally invasive surgery, enhancing intraoperative navigation, planning, and simulation through accurate modeling of dynamic scenes. Recent advances integrate dynamic radiance fields to address challenges like instrument occlusions and sparse viewpoints in endoscopic videos. However, achieving high fidelity in tissue deformation and topological variation, alongside real-time rendering for latency-sensitive applications, remains critical (Huang et al., 2024b; Liu et al., 2024b; Zhu et al., 2024c).
This task introduces distinct challenges compared to general dynamic scenes, including sparse training data from limited camera mobility in narrow cavities, tool occlusions obscuring critical regions, and single-view geometry ambiguities. Existing approaches mainly used additional depth guidance to infer the geometry of tissues (Huang et al., 2024b; Liu et al., 2024b; Zhu et al., 2024c). For instance, EndoGS (Zhu et al., 2024c) integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. EndoGaussian (Liu et al., 2024b) introduced two new strategies: holistic Gaussian initialization for dense initialization and spatiotemporal Gaussian tracking for modeling surface dynamics. Zhao et al. (Zhao et al., 2024d) argued that these methods suffer from under-reconstruction and proposed to alleviate this problem from frequency perspectives. In addition, EndoGSLAM (Wang et al., 2024c) and Gaussian Pancake (Bonilla et al., 2024) devised SLAM systems for endoscopic scenes and showed significant speed advantages.
@@ 5.6. Large-scale Scene Reconstruction
Large-scale scene reconstruction is a critical component in fields such as autonomous driving, aerial surveying, and AR/VR, demanding both photorealistic visual quality and real-time rendering capabilities. Before the emergence of 3D GS, the task has been approached using NeRF based methods, which, while effective for smaller scenes, often fall short in detail and rendering speed when scaled to larger areas (e.g., over 1.5 $km^{2}$). Though 3D GS has demonstrated considerable advantages over NeRFs, the direct application of 3D GS to large-scale environments introduces significant challenges. 3D GS requires an immense number of Gaussians to maintain visual quality over extensive areas, leading to prohibitive GPU memory demands and considerable computational burdens during rendering. For instance, a scene spanning 2.7 $km^{2}$ may require over 20 million Gaussians, pushing the limits of even the most advanced hardware (e.g., NVIDIA A100 with 40GB memory) (Liu et al., 2024a).
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training, a divide-and-conquer strategy (Liu et al., 2024a; Lin et al., 2024b; Ren et al., 2024; Kerbl et al., 2024) has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue (Cheng et al., 2024; Lin et al., 2024b). ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. LoD adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer (Lin et al., 2024b), or designing explicit LoD structures like the Octree (Ren et al., 2024) and hierarchy (Kerbl et al., 2024). Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process (Wu et al., 2024g, b; Xiong et al., 2024).
One prominent challenge lies in handling sparse or incomplete capture data, which can be mitigated through few-shot adaptation schemes (see Sec. 4.1) or generalizable priors (see “learning physical priors from large-scale data” in Sec. 7). Meanwhile, memory and computational bottlenecks can be addressed via distributed learning strategies (Zhao et al., 2024c), such as parameter partitioning across GPU clusters and parallel batched multi-view optimization.
@@ 5.7. Physics
The simulation of complex real-world dynamics, such as seed dispersal or fluid motion, is pivotal for applications spanning VR, animation, and scientific modeling, where realism hinges on accurate physical behavior. Advances in diffusion models have driven progress in 4D content generation, yet these methods might produce visually plausible results that violate fundamental physical laws. 3D GS emerges as a promising solution by embedding physical constraints and properties into scene representations, enabling both visually convincing and physically coherent simulations.
Existing methods differ in how they formulate and integrate physics-based priors into their frameworks. The most common approach is employing physics simulation engines (e.g., MLS-MPM (Hu et al., 2018)) to guide the dynamics generation. The material point method (Hu et al., 2018) and position based dynamics (Müller et al., 2007) — numerical methods used in computer graphics for simulating deformations in materials like fluids, granular media, and fracturing solids — have been extensively explored by the community through various customizations (Xie et al., 2024; Liu et al., 2024e; Borycki et al., 2024; Qiu et al., 2024; Huang et al., 2025; Zhang et al., 2024h; Jiang et al., 2024c; Feng et al., 2024a). Analytical material models, such as mass-spring systems, have also demonstrated success in approximating deformations by explicitly encoding material properties into 3D Gaussians (Zhong et al., 2024). Across these methods, 3D Gaussians are treated as discrete particles (with one exception (Shao et al., 2024a) using a continuous representation) and serve as computational units within the chosen simulator. Unknown material properties or physical parameters are typically learned through video-based supervision from conditional generative models.

### v9

@@ 5.4. Avatar
3D GS has emerged as a powerful tool for avatar reconstruction, primarily advancing along two directions: full-body modeling and head-centric modeling. For full-body avatars (Jena et al., 2023; Ye et al., 2023; Li et al., 2024j; Lei et al., 2024a; Moreau et al., 2024; Kocabas et al., 2024; Abdal et al., 2024; Zheng et al., 2024a; Hu and Liu, 2024; Jiang et al., 2024c; Hu et al., 2024b; Pang et al., 2024; Qian et al., 2024b; Yuan et al., 2024; Jung et al., 2023; Li et al., 2023a, 2024f; Zhai et al., 2025), the current methods typically anchor 3D Gaussians in a canonical space and deform them via parametric body models or cage-based rigging to model dynamic motions. These approaches adopt a hybrid deformation strategy: linear blend skinning handles rigid skeletal transformations such as joint rotations, while pose-conditioned deformation fields account for secondary non-rigid effects like muscle jiggles. For head avatars (Qian et al., 2024a; Dhamo et al., 2023; Xiang et al., 2023; Saito et al., 2024; Chen et al., 2023b; Zhao et al., 2024d; Zhou et al., 2024e; Li et al., 2024d; Rivero et al., 2024), the emphasis shifts to modeling intricate facial expressions, fine-grained geometry (e.g., wrinkles, hair (Luo et al., 2024)), and dynamic speech-driven animations. Techniques mainly combine parametric morphable face models (e.g., FLAME) with deformable 3D Gaussians, employing diffusion strategies and expression-aware deformation fields to disentangle rigid head poses from non-rigid facial movements. Both directions exploit the speed advantage and editability of GS to enable fast training, real-time rendering, and precise control over deformations, while addressing challenges in cross-frame correspondence, topology flexibility, and multi-view consistency.
@@ 5.5. Endoscopic Scene Reconstruction
Endoscopic Scene reconstruction is pivotal in robot-assisted minimally invasive surgery, enhancing intraoperative navigation, planning, and simulation. Recent advances integrate dynamic radiance fields to address challenges like instrument occlusions and sparse viewpoints in endoscopic videos. However, achieving high fidelity in tissue deformation and topological variation, alongside real-time rendering for latency-sensitive applications, remains critical (Huang et al., 2024e; Liu et al., 2024e; Zhu et al., 2024a).
This task introduces distinct challenges compared to general dynamic scenes, including sparse training data from limited camera mobility in narrow cavities, tool occlusions obscuring critical regions, and single-view geometry ambiguities. Existing approaches mainly used additional depth guidance to infer the geometry of tissues (Huang et al., 2024e; Liu et al., 2024e; Zhu et al., 2024a). For instance, EndoGS (Zhu et al., 2024a) integrated depth-guided supervision with spatial-temporal weight masks and surface-aligned regularization terms to enhance the quality and speed of 3D tissue rendering while addressing tool occlusion. EndoGaussian (Liu et al., 2024e) introduced two new strategies: holistic Gaussian initialization for dense initialization and spatiotemporal Gaussian tracking for modeling surface dynamics. Zhao et al. (Zhao et al., 2024a) argued that these methods suffer from under-reconstruction and proposed to alleviate this problem from frequency perspectives. In addition, EndoGSLAM (Wang et al., 2024b) and Gaussian Pancake (Bonilla et al., 2024) devised SLAM systems for endoscopic scenes and showed significant speed advantages.
@@ 5.6. Large-scale Scene Reconstruction
Large-scale scene reconstruction is a critical component in fields such as autonomous driving and aerial surveying. Before the emergence of 3D GS, the task has been approached using NeRF based methods, which, while effective for smaller scenes, often fall short in detail and rendering speed when scaled to larger areas (e.g., over 1.5 $km^{2}$). Though GS has demonstrated considerable advantages over NeRFs, the direct application of GS to large-scale environments introduces significant challenges. GS requires an immense number of Gaussians to maintain visual quality, leading to prohibitive GPU memory demands and computational burdens. For instance, a scene spanning 2.7 $km^{2}$ may require over 20 million Gaussians, pushing the limits of commonly used GPUs (e.g., NVIDIA A100 40GB) (Liu et al., 2024d).
To address the highlighted challenges, researchers have made significant strides in two key areas: i) For training, a divide-and-conquer strategy (Liu et al., 2024d; Lin et al., 2024a; Ren et al., 2024b; Kerbl et al., 2024) has been adopted, which segments a large scene into multiple, independent cells. This facilitates parallel optimization for expansive environments. An additional challenge lies in maintaining visual quality, as large-scale scenes often feature texture-less surfaces that can hamper the effectiveness of optimization such as Gaussian initialization and density control (Sec. 3.2). Enhancing the optimization algorithm presents a viable solution to mitigate this issue (Cheng et al., 2024; Lin et al., 2024a). ii) Regarding rendering, the adoption of the Level of Details (LoD) technique from computer graphics has proven instrumental. LoD adjusts the complexity of 3D scenes to balance visual quality with computational efficiency. Current implementations involve feeding only the essential Gaussians to the rasterizer (Lin et al., 2024a), or designing explicit LoD structures like the Octree (Ren et al., 2024b) and hierarchy (Kerbl et al., 2024). Furthermore, integrating extra input modalities like LiDAR can further enhanced the reconstruction process (Wu et al., 2024d, a; Xiong et al., 2024).
One prominent challenge lies in handling sparse or incomplete capture data, which can be mitigated through few-shot adaptation schemes (see Sec. 4.1) or generalizable priors (see “learning physical priors from large-scale data” in Sec. 7). Meanwhile, memory and computational bottlenecks can be addressed via distributed learning strategies (Zhao et al., 2024b), such as parameter partitioning across GPU clusters and parallel batched multi-view optimization.
@@ 5.7. Physics
The simulation of complex real-world dynamics, such as seed dispersal, fracture, or fluid motion, is pivotal for applications spanning VR, animation, scientific modeling, and emerging “world models”, where usefulness hinges not only on visual plausibility but also on adherence to physical laws. While recent diffusion-based 4D generation frameworks can synthesize visually compelling dynamics, they often optimize for appearance metrics and may violate basic conservation principles or material constraints. In this context, 3D GS provides an attractive bridge between graphics and simulation: Gaussians can be interpreted simultaneously as rendering primitives and as particles or elements in a numerical solver, making it natural to embed physical priors directly into the scene representation.
Existing methods differ mainly in how they couple Gaussians with explicit simulators. One line of work adopts classical schemes such as the material point method (MPM) (Hu et al., 2018) and position-based dynamics (PBD) (Müller et al., 2007), and replaces traditional grid or mesh discretizations with 3D Gaussians. In PhysGaussian (Xie et al., 2024), for instance, the same Gaussian set is used for both continuum-mechanics-based time integration and for splatting, so physical states evolve on exactly what is rendered. PBD-style pipelines such as Gaussian Splashing (Feng et al., 2024b) treat Gaussians as particles subject to constraints, enabling two-way coupling between deformable solids and fluids, while other systems like GASP (Borycki et al., 2024) derive mesh-like parameterizations from Gaussians to interface with conventional solvers. These approaches achieve high-fidelity dynamics within their target regimes but inherit the computational cost, stability issues, and parameter tuning burden of the underlying simulators. A complementary direction focuses on learning material parameters or effective constitutive models from data while keeping Gaussians as the canonical representation. Physics3D (Liu et al., 2024a) builds a viscoelastic MPM on top of 3DGS and estimates elastic and plastic properties by distilling priors from video diffusion models. DreamPhysics (Huang et al., 2025a), PhysDreamer (Zhang et al., 2024j), PhysMotion (Jiang et al., 2024a), and Feature Splatting (Qiu et al., 2024) similarly use generative video models as weak teachers for dynamics or material behavior, whereas mass-spring formulations such as Spring-Gaus (Zhong et al., 2024) encode stiffness and damping explicitly via springs attached to Gaussians. Most frameworks treat Gaussians as discrete particles inside differentiable simulators (with one exception of continuous formulations such as GausSim (Shao et al., 2024b)), and backpropagate through simulation and rendering to jointly adjust appearance and physics.
Despite rapid progress, physics-based GS is still far from a general-purpose physical world model. Current systems are often specialized to narrow regimes (e.g., a limited set of elastic behaviors), and extending a single Gaussian-based formulation to simultaneously handle rigid, elastic, plastic, granular, and fluid dynamics remains open. Interactions are typically restricted to a few objects in controlled scenes; robust multi-contact, multi-material interactions in cluttered environments still rely on manual priors, and long-horizon rollouts accumulate numerical and modeling errors. Moreover, physical supervision is weak: diffusion-based teachers target perceptual realism rather than quantitative accuracy, and existing datasets rarely provide ground-truth material parameters or stress/strain measurements. Bridging these gaps will require tighter integration between adaptive physics engines and Gaussian representations, better benchmarks for physical correctness, and learning schemes that combine explicit simulation with large-scale data-driven priors.
6. Performance Comparison
In this section, we provide empirical evidence by presenting the performance of several 3D GS algorithms previously discussed. The diverse applications of 3D GS across numerous tasks, coupled with the custom-tailored algorithmic designs for each task, render a uniform comparison of all 3D GS algorithms across a single task or dataset impracticable. For comprehensiveness, we provide a collection of representative datasets in Table Supplementary Material Following the taxonomy outlined Fig. 1. Due to the limited space, we have chosen several representative tasks for an in-depth performance evaluation. The performance scores are primarily sourced from the original papers, except where indicated otherwise.
Quantitative results (e.g., PSNR) can vary noticeably across implementations and hyperparameters (e.g., densification schedules and pruning resets), even on the same dataset. Since exhaustively enumerating these settings is beyond the scope of this survey, we link the original codebases and configurations in our benchmark repository to support reproducibility. Likewise, memory and storage footprint are important practical factors but are highly benchmark-dependent, varying with the dataset/scene (e.g., scale, duration, texture complexity, and number of views), the target quality, and implementation choices. Consequently, GS methods may exhibit either smaller or larger VRAM/storage than NeRF baselines depending on the dataset and the chosen capacity. We therefore provide the same repository to help readers reproduce and inspect footprint-quality trade-offs under specific datasets/settings of interest.
6.1. Performance Benchmarking: Localization
The localization task in SLAM involves determining the precise position and orientation of a robot or device.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent 3D GS based algorithms (Yugay et al., 2023; Matsuki et al., 2024; Keetha et al., 2024; Yan et al., 2024a) and six typical SLAM methods (Zhu et al., 2022; Sandström et al., 2023).
$\bullet$ Result: As shown in Table Supplementary Material, the recent 3D Gaussians based localization algorithms have a clear advantage over existing NeRF based dense visual SLAM. For example, SplaTAM (Keetha et al., 2024) achieves a trajectory error improvement of $\sim$50%, decreasing it from 0.52cm to 0.36cm compared to the previous state-of-the-art (SOTA) (Sandström et al., 2023). We attribute this to the dense and accurate 3D Gaussians reconstructed for scenes, which can handle the noise of real sensors. This reveals that effective scene representations can improve the accuracy of localization tasks.
6.2. Performance Benchmarking: Static Scenes
$\bullet$ Dataset: Replica (Straub et al., 2019) (Sec. 6.1) is used for comparison. The testing views are the same as (Sucar et al., 2021).
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent papers which introduce 3D Gaussians into their systems (Yugay et al., 2023; Matsuki et al., 2024; Keetha et al., 2024; Yan et al., 2024a), as well as three dense SLAM methods (Zhu et al., 2022; Sandström et al., 2023).
$\bullet$ Result: Table Supplementary Material shows that 3D Gaussians based systems generally outperform the three dense SLAM competitors. For example, Gaussian-SLAM (Yugay et al., 2023) establishes new SOTA and outperforms previous methods by a large margin. Compared to Point-SLAM (Sandström et al., 2023), GSSLAM (Matsuki et al., 2024) is about 578 times faster in achieving very competitive accuracy. In contrast to previous method (Sandström et al., 2023) that relies on depth information, such as depth-guided ray sampling, for synthesizing novel views, 3D GS based system eliminates this need, allowing for high-fidelity rendering for any views.
6.3. Performance Benchmarking: Dynamic Scenes
$\bullet$ Dataset: D-NeRF (Pumarola et al., 2021) dataset includes videos with 50 to 200 frames each, captured from unique viewpoints. It features synthetic, animated objects in complex scenes, with non-Lambertian materials. The dataset provides 50 to 200 training images and 20 test images per scene. The testing views are the same as the original paper (Pumarola et al., 2021).
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve five recent papers that model dynamic scenes with 3D GS (Yu et al., 2024a; Wu et al., 2024b; Lu et al., 2024c; Yang et al., 2024d, b), as well as six NeRF based approaches (Pumarola et al., 2021; Guo et al., 2023; Wang et al., 2023a).
$\bullet$ Result: From Table Supplementary Material we can observe that 3D GS based methods outperform existing SOTAs by a clear margin. The static version of 3D GS (Kerbl et al., 2023) fails to reconstruct dynamic scenes, resulting in a sharp drop in performance. By modeling the dynamics, D-3DGS (Yang et al., 2024d) outperforms the SOTA method, FFDNeRF (Guo et al., 2023), by 6.83dB in terms of PSNR. These results indicate the effectiveness of introducing additional properties or structured information to model the deformation of Gaussians so as to model the scene dynamics.
6.4. Performance Benchmarking: Human Avatar
$\bullet$ Result: Table Supplementary Material presents the numerical results of top-leading solutions in human avatar modeling. We observe that introducing 3D GS into the framework leads to consistent performance improvements in both rendering quality and speed. For instance, GART (Lei et al., 2024a) outperforms current SOTA, Instant-NVR (Geng et al., 2023), by 1.21dB in terms of PSNR. Considering the enhanced fidelity, inference speed and editability, 3D GS based avatar modeling may revolutionize the field of 3D animation, interactive gaming, etc.
6.5. Performance Benchmarking: Surgical Scenes
3D reconstruction from endoscopic video is critical to robotic-assisted minimally invasive surgery.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve three works which reconstruct dynamic 3D endoscopic scenes with GS (Huang et al., 2024e; Liu et al., 2024e; Zhao et al., 2024a), as well as three NeRF-based surgical reconstruction approaches (Wang et al., 2022; Zha et al., 2023).
$\bullet$ Result: Table Supplementary Material shows that introducing the explicit representation of 3D Gaussians leads to several significant improvements. For instance, EndoGaussian (Liu et al., 2024e) outperforms a strong baseline, EndoSurf (Zha et al., 2023), among all metrics. In particular, EndoGaussian demonstrates an approximate 200-fold increase in speed while consuming just 10% of the GPU resources, a gap that mainly stems from the GS pipeline and lightweight dynamic modeling rather than unusually heavyweight NeRF baselines. These impressive results attest to the efficiency of GS-based methods, which not only expedite processing but also minimize GPU load, thus easing the demands on hardware. Such attributes are significant for surgical application deployment, where optimized resource usage can be a key determinant of practical utility.
7. Future Research Directions
Recent Trends and The Big Picture. Recent work on GS has pushed it from a fast alternative to NeRFs towards a general representation for 3D world modeling. Early work showed that explicit 3D Gaussians can be optimized to capture complex static and dynamic scenes with high fidelity, inspiring follow-up research on more compact, generalizable, and semantics-aware splat-based models (Lee et al., 2024b; Liu et al., 2024b; Gao et al., 2025). As an explicit scene representation, GS has been applied across a wide range of vision, graphics, and robotics applications (Tang et al., 2024b; Abou-Chakra et al., 2023; Feng et al., 2025; Abou-Chakra et al., 2024). A major recent trend is the shift from per-scene optimization to feed-forward Gaussian prediction. Early pioneering works showed that 3D Gaussians can be inferred directly from sparse views using lightweight, pixel-level feature extractors (e.g., U-Nets) (Charatan et al., 2024; Szymanowicz et al., 2024b). To improve scalability from object-centric reconstructions to complex scenes, more recent approaches have adopted Transformer-based models (Zhang et al., 2024f; Xu et al., 2024b). This architectural shift has naturally driven efforts to process increasingly large input sets; by innovating on sequence processing – such as integrating hybrid attention mechanisms (Ziwen et al., 2025b, a) or test-time training (Zhang et al., 2026; Wang et al., 2026) – current models can now ingest massive multi-view contexts for wide-coverage, scene-level modeling. Simultaneously, this feed-forward paradigm is rapidly expanding its functional scope: recent frameworks can now jointly infer camera parameters from uncalibrated image collections (Jiang et al., 2025), extract robust geometric priors to stabilize downstream prediction tasks (Wang et al., 2025), and extend spatial prediction into the temporal domain for 4D dynamic generation (Ren et al., 2024a; Ma et al., 2025). Together, these trends point towards Gaussian-based world models that support not only high-quality rendering and reconstruction but also interactive perception, reasoning, and action, with demonstrations in robotics and embodied AI. Concrete discussions on some important directions are given in the following.
$\bullet$ Physics- and Semantics-aware Scene Representation. As a new, explicit scene representation technique, 3D Gaussian offers transformative potential beyond merely enhancing novel-view synthesis. It has the potential to pave the way for simultaneous advancements in scene reconstruction and understanding by devising physics- and semantics-aware 3D GS systems. While significant progress has been made in physics (Sec. 5.7) and semantics (Wang et al., 2024a; Qu et al., 2024; Ji et al., 2024d; Liao et al., 2024; Choi et al., 2024; Ji et al., 2024b) individually, there remains considerable untapped potential in their synergistic integration. This is poised to revolutionize a range of fields and downstream applications. For instance, incorporating prior knowledge such as the general shape of objects can reduce the need for extensive training viewpoints (Szymanowicz et al., 2024b; Charatan et al., 2024) while improving geometry/surface reconstruction (Guédon and Lepetit, 2024; Li et al., 2024h). A critical metric for assessing scene representation is the quality of its generated scenes, which encompasses challenges in geometry, texture, and lighting fidelity (Huang et al., 2024d; Chen et al., 2024c; Gao et al., 2023). By merging physical principles and semantic information within the 3D GS framework, one can expect that the quality will be enhanced, thereby facilitating dynamics modeling (Xie et al., 2024; Jiang et al., 2024a), editing (Cen et al., 2023; Zhou et al., 2024b), generation (Tang et al., 2024b; Yi et al., 2024), and beyond. In a nutshell, pursuing this advanced and versatile scene representation opens up new possibilities for innovation in computational creativity and practical applications across diverse domains.
$\bullet$ Learning Physical Priors from Large-scale Data. As we explore the potential of physics- and semantics-aware scene representations, leveraging large-scale datasets to learn generalizable, physical priors emerges as a promising direction. The goal is to model the inherent physical properties and dynamics embedded within real-world data, transforming them into actionable insights that can be applied across various domains such as robotics and visual effects. Establishing a learning framework for extracting these generalizable priors enables the application of these insights to specific tasks in a few-shot manner. For instance, it allows for rapid adaptation to new objects and environments with minimal data input. Furthermore, integrating physical priors can enhance not only the accuracy and quality of generated scenes but also their interactive and dynamic qualities. This is particularly valuable in AR/VR environments, where users interact with virtual objects that behave in ways consistent with their real-world counterparts. However, the existing body of work on capturing and distilling physics-based knowledge from extensive 2D and 3D datasets remains sparse. Notable efforts in related area include the continuum mechanics based GS systems (Sec. 5.7), and the generalizable Gaussian representation based on multi-view stereo (Liu et al., 2024b). Further exploration on real2sim and sim2real might offer viable routes for advancements in this field.
$\bullet$ Modeling Internal Structures of Objects with 3D GS. Despite the ability of 3D GS to produce highly photorealistic renderings, modeling internal structures of objects (e.g., for a scanned object in computed tomography) within the current GS framework presents a notable challenge. Due to the splatting and density control process, the current representation of 3D Gaussian is unorganized and cannot align well with the object’s actual internal structures. Moreover, there is a strong preference in various applications to depict objects as volumes (e.g., computed tomography). However, the disordered nature of 3D GS makes volume modeling particularly difficult. Li et al. (Li et al., 2023b) used 3D Gaussians with density control as the basis for the volumetric representation and did not involve the splatting process. X-Gaussian (Cai et al., 2024a) involves the splatting process for fast training and inference but cannot generate volumetric representation. Using 3D GS to model the internal structures of objects remains unanswered and deserves further exploration.
$\bullet$ 3D GS for Simulation in Autonomous Driving and beyond. Collecting real-world datasets for autonomous driving is both expensive and logistically challenging, yet crucial for training effective image-based perception systems. To mitigate these issues, simulation emerges as a cost-effective alternative, enabling the generation of synthetic datasets across diverse environments. However, the development of simulators capable of producing photorealistic and diverse synthetic data is fraught with challenges. These include achieving a high level of quality, accommodating various control methods, and accurately simulating a range of lighting conditions. While early efforts (Zhou et al., 2024d; Yan et al., 2024b; Zhou et al., 2024a) in reconstructing urban/street scenes with 3D GS have been encouraging, they are just the tip of the iceberg in terms of the full capabilities. There remain numerous critical aspects to be explored, such as the integration of user-defined object models, the modeling of physics-aware scene changes (e.g., the rotation of vehicle wheels), and the enhancement of controllability and quality (e.g., in varying lighting conditions). Mastery of these capabilities would not only advance autonomous systems but also redefine computational understanding of physical spaces — a leap with implications for world models, spatial intelligence, embodied AI, and beyond.
$\bullet$ Empowering 3D GS with More Possibilities. Despite the significant potential of 3D GS, the full scope of applications for 3D GS remains largely untapped. A promising avenue for exploration involves augmenting 3D Gaussians with additional attributes (e.g., linguistic and spatiotemporal properties as mentioned in Sec. 4.5) and introducing structured information (e.g., spatial MLPs and grids as mentioned in Sec. 4.6), tailored for specific applications. Moreover, recent studies have begun to unveil the capability of 3D GS in several domains, e.g., point cloud registration (Chang et al., 2024), image representation and compression (Zhang et al., 2024k), and fluid synthesis (Feng et al., 2024b). These findings highlight a significant opportunity for interdisciplinary scholars to explore 3D GS further.
8. Conclusions
Supplementary Material
Due to the page limit of the main manuscript, we provide additional figures and tables in this supplementary material. Specifically, we include i) a visualization of the 3D GS growth process; ii) a visual comparison between NeRFs and 3D GS; iii) a demonstration of tile-based parallel rendering; iv) a consolidated summary table of the commonly used datasets; and v) detailed quantitative performance-comparison tables. These figures and tables are referenced from the main text and are intended to improve completeness without interrupting the main narrative.
Figure S1. The number of published papers and GitHub stars on 3D GS. Statistics sourced from # Papers and # GitHub Stars.
$x,y,z,\theta,\phi$
$c\penalty 10000\ \&\penalty 10000\ \sigma$
Figure S2. NeRFs vs. 3D GS. (a) NeRF samples along the ray and then queries the MLP to obtain colors and densities, which can be seen as a backward mapping (ray tracing). (b) 3D GS projects all 3D Gaussians into the image space (i.e., splatting) and then performs parallel rendering, which can be viewed as a forward mapping (rasterization).
5
Figure S3. An illustration of the tile based parallel (at the pixel-level) rendering. All the pixels within a tile (Tile1 here) access the same ordered Gaussian list stored in a shared memory for rendering. As the system processes each Gaussian sequentially, every pixel in the tile evaluates the Gaussian’s contribution according to the distance (i.e., the $\exp$ term in Eq. 6). Therefore, the rendering for a tile can be completed by iterating through the list of Gaussians just once. The computation for the red Gaussian follows a similar way and is omitted here for simplicity.
Table S1. Summary of major 3DGS research directions reviewed in Section 4. For each direction, representative works are listed in chronological order, providing a rough timeline of progress.
  Direction Key ideas Representative works Advantages Limitations / challenges
§4.1 3DGS for Sparse Input Handle few-view / sparse inputs via regularization or learned priors (Xiong et al., 2023; Zhu et al., 2024c; Charatan et al., 2024; Szymanowicz et al., 2024b; Li et al., 2024c; Swann et al., 2024; Chen et al., 2024d; Wewer et al., 2024; Xu et al., 2024b; Shen et al., 2024; Zhang et al., 2024e) Supports limited viewpoints; enables single-image or few-shot NVS; leverages strong 3D priors Trade-off between geometry quality and speed; sensitive to prior/regularization choice; dynamic scenes remain difficult
§4.2 Memory-efficient 3DGS Reduce Gaussian count and/or compress Gaussian attributes for compact storage (Fan et al., 2023; Navaneet et al., 2023; Lee et al., 2024b; Morgenstern et al., 2023; Zhang et al., 2024k; Niedermayr et al., 2024; Chen et al., 2024b; Papantonakis et al., 2024; Fang and Wang, 2024) Lower memory footprint; scalable to large scenes and datasets; facilitates deployment and streaming Quality drop at high compression; training-time memory is still high; designing reusable codebooks is open
§4.3 Photorealistic 3D GS Improve aliasing, reflection, blur, and shading to boost visual fidelity (Yu et al., 2024c; Yan et al., 2024c; Jiang et al., 2024b; Lee et al., 2024a; Bolanos et al., 2024; Radl et al., 2024; Yang et al., 2024c; Zhao et al., 2024c; Dahmani et al., 2024; Li et al., 2024h; Liang et al., 2024b; Seiskari et al., 2024; Song et al., 2024) Sharper details; more realistic materials and lighting; increased robustness on real captures Pipelines become more complex; many methods target specific artifacts; may increase training and rendering cost
§4.4 Improved Optimization Algorithms Better optimization and regularization pipelines for more stable training (Fu et al., 2024; Lu et al., 2024b; Jung et al., 2024; Li et al., 2024h; Yu et al., 2024b; Zhang et al., 2024d; Huang et al., 2024c; Li et al., 2024e) Faster convergence; fewer floaters/artifacts; more robust reconstruction in challenging scenes Still per-scene optimization; hyperparameter and schedule tuning; limited work on meta-learning or few-shot GS
§4.5 3D Gauss-ians with More Properties Attach semantic or spatiotemporal features to Gaussians for richer scene understanding (Shi et al., 2024; Qin et al., 2024; Zhou et al., 2024b; Ye et al., 2024; Cen et al., 2023; Yang et al., 2024b; Zuo et al., 2024) Enables open-vocabulary querying, semantic editing, spatiotemporal modeling, etc. Higher memory and computation; depends on large-scale supervision; Evaluation protocols are still emerging
§4.6 Hybrid Re-presentations Combine Gaussians with structured modules (MLPs, deformation fields) for specific tasks (Yang et al., 2024d; Wu et al., 2024b; Xu et al., 2024c; Saroha et al., 2024) Adds structure and controllability; effective for avatars, dynamic scenes, stylization, etc. More complex architectures; less generic; partially loses the simplicity of pure GS
§4.7 New Rendering Algorithms Devise ray tracing / physically-based and improved rasterization for Gaussians (Moenne-Loccoz et al., 2024; Mai et al., 2024; Condor et al., 2025) Better handling of visibility, secondary rays, and complex cameras; improved physical correctness Higher computational cost; integration with real-time pipelines is still in early stages; APIs and tooling are immature
Table S2. Comparison of regularization-based and generalizability-based 3DGS methods for sparse input.
  Dimension Regularization-based Generalizability-based
Inference speed Slow; per-scene optimization Fast; feed-forward
Reconstruction quality High detail; strong geometry Lower accuracy; priors may hallucinate
Views required at test time Typically multi-view Few views; even single image
Robustness to unseen regions Limited; overfits observed views Depends on prior; plausible but inexact
Typical use cases High-fidelity reconstruction Real-time / interactive NVS
Representative works (Li et al., 2024c; Zhu et al., 2024c; Chen et al., 2024d; Zhang et al., 2024e) (Charatan et al., 2024; Szymanowicz et al., 2024b; Xu et al., 2024b; Szymanowicz et al., 2024a)